OpenStudy (anonymous):

Help!! Evaluate the expression exactly... Sin(-90) + Csc(90)

OpenStudy (anonymous):

sin(-90)=-1 and cos(90) = so its -1

OpenStudy (anonymous):

from the trigonometric circle

OpenStudy (campbell_st):

actually its Csc(90) which is 1 so easy from here

OpenStudy (anonymous):

csc>cosine ??

OpenStudy (campbell_st):

the reason csc(90) = 1 is that csc = 1/sin and sin(90) = 1 so csc(90) = 1/1

OpenStudy (anonymous):

not really familliary with the englisht tearm

OpenStudy (campbell_st):

cosecant... is csc... cosine is cos... cosecant = 1/sin

OpenStudy (anonymous):

my bad